# Glauber dynamics for the hard-core model on bounded-degree $H$-free graphs

### v1

@@ 1. Introduction
The complete bipartite graph $K_{3,1}$ is known as a claw. A graph is said to be claw-free if it does not contain a claw as an induced subgraph. The hard-core model is specified by a graph $G$ and a fugacity $\lambda>0$. These inputs define a probability distribution on independent sets of a graph, the so-called hard-core distribution. Higher values of $\lambda$ favour larger independent sets, leading to denser configurations. (Precise definitions will be given presently.)
@@
It is natural to wonder whether excluding graphs other than the claw can lead to optimal mixing time for all $\lambda$. There is an extensive literature on independent sets in graph families that exclude a fixed connected graph $H$ as an induced subgraph. (From now on we restrict attention to connected $H$.) In this work, subdivided claws play a crucial role [1, 10]. A subdivided claw is obtained from a claw by subdividing the three edges by arbitrary number of intermediate vertices. Alternatively, a subdivided claw is a tree with a single vertex of degree 3, with all other vertices having degree 1 or 2. For $1\leq i\leq j\leq k$, we write $S_{i,j,k}$ for the subdivided claw whose leaves are at distance $i$, $j$ and $k$ from the degree-3 vertex. The smaller subdivided claws have special names: $S_{1,1,1}$ (as we have seen) is the claw, $S_{1,1,2}$ the fork, $S_{1,2,2}$ is the E, and $S_{1.2.3}$ is the skew star. See Figure 1.
@@
It soon becomes apparent that the interesting cases for us are when the excluded graph $H$ is a subdivided claw. Informally, the reason is this: if $H$ is not a subdivided claw or a path, then it contains either a vertex of degree greater than 3, or a cycle, or two vertices of degree 3 connected by a path. We can exclude the first case by working with instance graphs $G$ with maximum degree 3, and the other two cases by working with graphs whose edges have been ‘stretched’ or subdivided by sufficiently many vertices. This intuition is formalised in Section 6. Note that excluding a path in a bounded-degree graph leads to a finite graph class. The (fairly routine) conclusion is that Glauber dynamics mixes in exponential time for sufficiently large $\lambda$ (depending on $\Delta$ and $H$ but not on $n=|V(G)|$) when $H$ is not a subdivided claw or a path.
@@ 2. Preliminaries
We think of $\mu_{G,\lambda}$ being a distribution on $2^{V(G)}$ with support $\mathcal{I}_{G}$. For a vertex $v$ in $G$ we write $\mu_{G,\lambda}^{(v,0)}$ (respectively $\mu_{G,\lambda}^{(v,1)}$) for the distribution on independent sets obtained from $\mu_{G,\lambda}$ by conditioning on $v$ being excluded (respectively included) in the independent set.
@@
    • with the remaining probability set $I^{\prime}\leftarrow I\setminus\{v\}$;.
@@
Then the mixing time of the Markov chain is defined to be $t_{\mathrm{mix}}=\min\{t:d(t)\leq\frac{1}{4}\}$. The constant $\frac{1}{4}$ is arbitrary, subject to lying in the interval $(0,\frac{1}{2})$. It is standard [16, §4.5] that convergence to stationary is exponential, when time is measured in units of $t_{\mathrm{mix}}$, specifically, $d(\ell\,t_{\mathrm{mix}})\leq 2^{-\ell}$, for all $\ell\in\mathbb{N}$. The central question is then: is $t_{\mathrm{mix}}$ bounded by a polynomial in $n=|V(G)|$ and, more particularly, whether $t_{\mathrm{mix}}=O(n\log n)$ (with the constant of proportionality being a function of $\lambda$ and $\Delta$).
@@
A simple way to see this is by monotone coupling of Markov chains [13, §3.10.3]. Modify the Markov chain from Figure 2 by removing the test $\Gamma_{G}(v)\cap I=\emptyset$ and the ‘else’ part of the conditional statement. Thus the probabilistic update is done whatever the state of the vertices adjacent to $v$. This modified process clearly converges to the product distribution $\beta_{G,\lambda}$. Now run the modified process starting from the all-1 configuration in parallel with Glauber dynamics starting from the all-0 configuration. Use the same random choice of vertex $v$ at each step, and optimally couple the updates. It is easy to see that the first process remains above the second in the product order. The upper process converges to $\beta_{G,\lambda}$ and the lower to $\mu_{G,\lambda}$. ∎
@@
$\frac{\lambda}{(1+\lambda)^{\Delta+1}}\leq\Pr(v\in I)\leq\frac{\lambda}{1+\lambda}.$
@@
The upper bound is immediate from Lemma 1. For the lower bound, note that $\Pr(I\cap\Gamma(v)=\emptyset)\geq 1/(1+\lambda)^{\Delta}$, and that $\Pr(v\in I\mid I\cap\Gamma(v)=\emptyset)\geq\lambda/(1+\lambda)$, again from Lemma 1. ∎
@@
For a graph $G$ and vertex $v\in V(G)$ let $\mu_{G,\lambda}^{(v,0)}$ (respectively, $\mu_{G,\lambda}^{(v,1)}$) denote the hard-core distribution on $G$ conditioned on vertex $v$ being excluded from (respectively, included in) the independent set. We will be using the powerful spectral independence machinery of Anari, Liu, Oveis Gharan and Vinzant [4] as sharpened by Cryan, Guo and Mousa [11] and Chen, Liu and Vigoda [8]. However, we can ease our task by using Chen and Gu’s packaging of the machinery for situations such as ours. Indeed, to show optimal mixing we just need to construct a coupling that achieves bounded Wasserstein distance $W_{1}\big{(}\mu_{G,\lambda}^{(v,0)},\mu_{G,\lambda}^{(v,1)}\big{)}$.
@@
Let $\mathcal{G}$ be a hereditary class of graphs with degree bound $\Delta$, and $\lambda$ be positive. Suppose there exists $\eta$ such that for all $G\in\mathcal{G}$ and all $v\in V(G)$ we have $W_{1}(\mu_{G,\lambda}^{(v,0)},\mu_{G,\lambda}^{(v,1)})\leq\eta$, where $\mu_{G,\lambda}$ is the hard-core distribution with fugacity $\lambda$ on $V(G)$. Then the mixing time of Glauber dynamics for the hard-core model on $G\in\mathcal{G}$ is bounded above by $O(n\log n)$, where $n=|V(G)|$. The constant implicit in the $O$-notation depends on $\lambda$, $\Delta$ and $\eta$.
@@
This follows from [7, Thm 9 and Lemma 10], [^1] The quoted results are stated for edge-based configurations rather than vertex-based, but this is inessential. which builds on earlier work, such as [8]. In order to make the correspondence we need to discuss briefly the notion of ‘pinning’. The distributions $\mu_{G,\lambda}^{(v,0)}$ and $\mu_{G,\lambda}^{(v,1)}$ are examples of distributions obtained by pinning, in this instance by pinning vertex $v$ to 0 and 1, respectively. More generally, we may consider pinning a set of vertices, say $\Lambda\subset V(G)$, to values given by $\tau:\Lambda\to\{0,1\}$, resulting in a marginal distribution $\mu_{G,\lambda}^{(\Lambda,\tau)}$ on $V(G)\setminus\Lambda$. Lemma 10 of [7] quantifies over pairs of pinnings $(\Lambda,\tau)$ and $(\Lambda,\tau^{\prime})$ where $\tau$ and $\tau^{\prime}$ differ at exactly one point. However, the hard-core distribution has a pleasant property: pinning a vertex $v$ to 0 is equivalent to removing $v$ from the graph, while pinning $v$ to 1 is equivalent to removing $v$ and all its neighbours. So, in the case of the hard-core model we need only quantify over single vertex pinnings, since we are restricting attention to graph classes which are closed under taking induced subgraphs.
From our premisses, Lemma 10 of [7] tells us that that $\mu_{G,\lambda}$ is ‘$\eta$-spectrally bounded’. For our purposes we don’t need to define this concept as we are merely going to feed this information about $\mu_{G,\lambda}$ into Theorem 9 of [7]. The remaining premiss required for this theorem is that $\mu_{G,\lambda}$ is ‘$b$-marginally bounded’. This just means that $\Pr(\sigma(v)=1)$ is bounded away from both 0 and 1 by a margin $b$. But this is exactly what is guaranteed by Corollary 2, with $b=\min\{\lambda/(1+\lambda)^{\Delta+1},1/(1+\lambda)\}$. ∎
@@ 3. $S_{1,1,t}$-free graphs (claw-free and upwards)
Suppose $\Delta\geq 3$ is a degree bound and $H$ a fixed graph. Let $G$ be an $H$-free connected graph of maximum degree $\Delta$. Chen and Gu [7, Thm 4] show optimal mixing of Glauber dynamics for the hard-core model on $G$, for any given fugacity $\lambda$, when the excluded graph $H$ is the claw. In fact, optimal mixing also holds for $H=S_{1,1,t}$ with $t\geq 2$, for a rather dull reason. Define
@@
  • $A=\{\{v,w_{1}\}\}$: the set $\{v,a,b,w_{1},\ldots,w_{s-1}\}$ induces $S_{1,1,t}$;
  • $A=\{\{a,w_{1}\}\}$: the set $\{v,b,c,a,w_{1},\ldots,w_{s-2}\}$ induces $S_{1,1,t}$;
  • $A=\{\{v,w_{1}\},\{a,w_{1}\}\}$: the set $\{v,b,c,w_{1},\ldots,w_{s-1}\}$ induces $S_{1,1,t}$;
  • $A\supseteq\{\{a,w_{1}\},\{b,w_{1}\}\}$: the set $\{w_{1},a,b,w_{2},\ldots,w_{s}\}$ induces $S_{1,1,t}$.
@@
This is immediate from [7, Thm 4] and Corollary 5. ∎
@@ 4. $S_{1,2,t}$-free graphs (E-free and upwards)
Before heading off to larger graph classes it is instructive to go back to a precursor of [7]. Van den Berg and Brouwer [20] construct an explicit coupling for matchings in a bounded-degree graph that (in modern terms) bounds spectral independence. This almost immediately gives optimal mixing of Glauber dynamics for the monomer-dimer model at all temperatures. (Of course, van den Berg and Brouwer did not have the spectral independence technology, so they were only able to deduce low-degree polynomial mixing time.) Their method extends easily to independent sets in claw-free graphs (though this was not stated in their paper). In this section, we rederive the result of Chen and Gu, as the proof will provide a base from which we may progress to large subdivided claws.
  • Let $R$ is a random ‘red’ independent set from $\mu_{G,\lambda}^{(v,1)}$
@@
  • Consider the subgraph $G[R\cup B]$ of $G$ induced by the union of the two independent sets. Let $C\subseteq V(G)$ be the connected component of $G[R\cup Y]$ containing $v$. We refer to $C$ as a cluster.
@@
Let $G$ be a connected claw-free graph of maximum degree $\Delta$, $v\in V(G)$, and $\lambda>0$. Then the coupling of Figure 3 is correct: $R$ is a sample from $\mu_{G,\lambda}^{(v,1)}$ and $B$ is a sample from $\mu_{G,\lambda}^{(v,0)}$,
@@
  Cluster $\textsf{Cluster BFS}(v)$
@@
The claimed bound on $W_{1}$ is evidenced by the coupling just analysed. As $|R\oplus B|=|C|$, we just have to demonstrate that $|C|$ is small in expectation. Note that $G[C]$ is bipartite. It is straightforward to see that a connected bipartite claw-free graph can have no vertices of degree 3 or greater, so is either a path or a cycle. As paths and cycles are one-dimensional structures, we do not expect them to propagate far, and this is indeed the case. Imagine that we reveal the colours of vertices in a breadth-first fashion, starting with $v$ at time 0. The breadth-first search is presented as Algorithm 1. This is a little heavy for current purposes, but its generality will be useful later. As the cluster is either a path or a cycle, we have $|L_{j}|\leq 2$ for all levels $j$. Since the hard core distribution is dominated by the product of independent Bernoullli distributions (Lemma 1), on each iteration, with probability at least $p=(1/(\lambda+1))^{2\Delta}$, the breadth-first search will halt. So $\mathop{\mathbb{E\hbox{}}}|C|=\mathop{\mathbb{E\hbox{}}}[|R\oplus B|]$ is bounded above by the expectation of a random variable having a geometric distribution with success probability $(\lambda+1)^{-2\Delta}$, which is $(\lambda+1)^{2\Delta}$. ∎
@@
The first interesting case beyond claw free is E-free (i.e., $S_{1,2,2}$-free). Figure 4 illustrates an infinite family of E-free graphs which are not claw free. Indeed, by iterating the basic building block either horizontally or vertically, or both, we can appreciate that the family includes graphs that contain induced copies of $S_{1,1,t}$ for arbitrarily large $t$. (An instance of a fork is picked out in red in Figure 4.) So, by excluding E, we have escaped from the constraints of Corollary 5 into new, apparently interesting families of graphs.
@@
The cluster containing $v$ is either of bounded size or is the blow-up of a path or cycle. The layers $L_{i}$ created by the breadth-first search of Algorithm 1 are of bounded size, so the depth of the search is again bounded by a geometric random variable. ∎

### v2

@@ 1. Introduction
The complete bipartite graph $K_{1,3}$ is known as a claw. A graph is said to be claw-free if it does not contain a claw as an induced subgraph. The hard-core model is specified by a graph $G$ and a fugacity $\lambda>0$. These inputs define a probability distribution on independent sets of a graph, the so-called hard-core distribution. Higher values of $\lambda$ favour larger independent sets, leading to denser configurations. (Precise definitions will be given presently.)
@@
It is natural to wonder whether excluding graphs other than the claw can lead to optimal mixing time for all $\lambda$. There is an extensive literature on independent sets in graph families that exclude a fixed connected graph $H$ as an induced subgraph. (From now on we restrict attention to connected $H$.) In this line of work, subdivided claws play a crucial role [1, 13]. A subdivided claw is obtained from a claw by subdividing the three edges by an arbitrary number of intermediate vertices. Alternatively, a subdivided claw is a tree with a single vertex of degree 3, and all other vertices having degree 1 or 2. For $1\leq i\leq j\leq k$, we write $S_{i,j,k}$ for the subdivided claw whose leaves are at distance $i$, $j$ and $k$ from the degree-3 vertex. The smaller subdivided claws have special names: $S_{1,1,1}$ (as we have seen) is the claw, $S_{1,1,2}$ the fork, $S_{1,2,2}$ is the E, and $S_{1.2.3}$ is the skew star. See Figure 1.
@@
It soon becomes apparent that the interesting cases for us are when the excluded graph $H$ is a subdivided claw. Informally, the reason is this: if $H$ is not a subdivided claw or a path, then it contains either a vertex of degree greater than 3, or a cycle, or two vertices of degree 3 connected by a path. We can exclude the first case by working with instance graphs $G$ with maximum degree 3, and the other two cases by working with graphs whose edges have been ‘stretched’ or subdivided by sufficiently many vertices. This intuition is formalised in Section 6. Note that excluding a path in a bounded-degree graph leads to a finite graph class. The (fairly routine) conclusion, expressed in Theorem 18, is that Glauber dynamics mixes in exponential time for sufficiently large $\lambda$ (depending on $\Delta$ and $H$ but not on $n=|V(G)|$) when $H$ is not a subdivided claw or a path.
Our main result, Theorem 17 in Section 6, asserts that bounded-degree graphs that exclude any subdivided claw support optimal mixing. Together with the matching negative result just mentioned, we obtain the following dichotomy. If $\mathcal{H}$ is any collection of graphs, we say that a graph $G$ is $\mathcal{H}$-free if $G$ is $H$-free (does not contain $H$ as an induced subgraph) for all $H\in\mathcal{H}$.
Theorem 1.
Let $\mathcal{H}$ be any finite set of connected graphs, and $\Delta\geq 3$. Consider the hard-core model on $\mathcal{H}$-free graphs of maximum degree $\Delta$. If $\mathcal{H}$ contains a subdivided claw or a path then Glauber dynamics mixes in optimal $O(n\log n)$ time, at any fixed fugacity $\lambda>0$, where $n$ is the number of vertices of $G$. Otherwise, for some sufficiently large fixed $\lambda$ (depending on $\mathcal{H}$), the mixing time is exponential in $n$ in the worst case, even when $\Delta=3$.
The above statement is essentially a composite of Theorems 17 and 18, and the (very short) proof can be found in Section 6.
@@
It is interesting to note that claw-free graphs arise in nature as crystal structures, for example the kagome and pyrochlore lattices. However, I am unaware of any real-life crystals that are free of subdivided claws without in fact being claw-free.
In related work, Chudnovsky and Seymour [14] showed that the partition function of the hard-core model on claw-free graphs has only (negative) real roots. The existence of a zero-free region that includes the positive real axis has implications for the existence of deterministic algorithms for approximating the partition function, through work of Barvinok [6] and Patel and Regis [24]. For fork-free graphs, a zero-free region, smaller but still including the positive real axis, has been identified by Bencs [8].
@@
Finally, although we concentrate on the hard-core model here, a result of a similar flavour has been obtained for the antiferromagnetic Ising model. For this model, Glauber dynamics has been shown to have optimal mixing time on the class of line graphs at any non-zero temperature however small (i.e., for any interaction strength however large). This result is due to Chen, Liu and Vigoda [11], strengthening the mixing time bound of Dyer, Heinrich, Jerrum and Müller [16] from polynomial to optimal.
@@ 2. Preliminaries
We think of $\mu_{G,\lambda}$ being a distribution on $2^{V(G)}$ with support $\mathcal{I}_{G}$.
@@
    • with the remaining probability set $I^{\prime}\leftarrow I\setminus\{v\}$;
@@
Then the mixing time of the Markov chain is defined to be $t_{\mathrm{mix}}=\min\{t:d(t)\leq\frac{1}{4}\}$. The constant $\frac{1}{4}$ is arbitrary, subject to lying in the interval $(0,\frac{1}{2})$. It is standard [21, §4.5] that convergence to stationary is exponential, when time is measured in units of $t_{\mathrm{mix}}$, specifically, $d(\ell\,t_{\mathrm{mix}})\leq 2^{-\ell}$, for all $\ell\in\mathbb{N}$. The central question is then: is $t_{\mathrm{mix}}$ bounded by a polynomial in $n=|V(G)|$ and, more particularly, is $t_{\mathrm{mix}}=O(n\log n)$ (with the constant of proportionality being a function of $\lambda$ and $\Delta$)?
@@
A simple way to see this is by monotone coupling of Markov chains [18, §3.10.3]. Modify the Markov chain from Figure 2 by removing the test $\Gamma_{G}(v)\cap I=\emptyset$ and the ‘else’ part of the conditional statement. Thus the probabilistic update is done whatever the state of the vertices adjacent to $v$. This modified process clearly converges to the product distribution $\beta_{G,\lambda}$. Now run the modified process starting from the all-1 configuration in parallel with Glauber dynamics starting from the all-0 configuration. Use the same random choice of vertex $v$ at each step, and optimally couple the updates. It is easy to see that the first process remains above the second in the product order. The upper process converges to $\beta_{G,\lambda}$ and the lower to $\mu_{G,\lambda}$. ∎
@@
$\frac{\lambda}{(1+\lambda)^{\Delta+1}}\leq\mathop{\mathbb{P\hbox{}}}(v\in I)\leq\frac{\lambda}{1+\lambda}.$
@@
The upper bound is immediate from Lemma 2. For the lower bound, note that $\mathop{\mathbb{P\hbox{}}}(I\cap\Gamma(v)=\emptyset)\geq 1/(1+\lambda)^{\Delta}$, again from Lemma 2, and that $\mathop{\mathbb{P\hbox{}}}(v\in I\mid I\cap\Gamma(v)=\emptyset)\geq\lambda/(1+\lambda)$.∎
@@
For a graph $G$ and vertex $v\in V(G)$ let $\mu_{G,\lambda}^{(v,0)}$ (respectively, $\mu_{G,\lambda}^{(v,1)}$) denote the hard-core distribution on $G$ at fugacity $\lambda$ conditioned on vertex $v$ being excluded from (respectively, included in) the independent set. We will be using the powerful spectral independence machinery of Anari, Liu, Oveis Gharan and Vinzant [5] and Anari, Liu and Oveis Gharan [4], as sharpened by Cryan, Guo and Mousa [15] and Chen, Liu and Vigoda [10]. However, we can ease our task by using Chen and Gu’s packaging of the machinery for situations such as ours. Indeed, as we shall see presently, to show optimal mixing we just need to construct a coupling that achieves bounded Wasserstein distance $W_{1}\big{(}\mu_{G,\lambda}^{(v,0)},\mu_{G,\lambda}^{(v,1)}\big{)}$.
A graph class is said to be hereditary if it is closed under taking induced subgraphs. Note that graph classes defined by excluding induced subgraphs, with or without a degree bound, are hereditary.
@@
Let $\mathcal{G}$ be a hereditary class of graphs with degree bound $\Delta$, and $\lambda$ be positive. Suppose there exists $\eta>0$ such that for all $G\in\mathcal{G}$ and all $v\in V(G)$ we have $W_{1}(\mu_{G,\lambda}^{(v,0)},\mu_{G,\lambda}^{(v,1)})\leq\eta$, where $\mu_{G,\lambda}$ is the hard-core distribution with fugacity $\lambda$ on $V(G)$. Then the mixing time of Glauber dynamics for the hard-core model on $G\in\mathcal{G}$ at fugacity $\lambda$ is bounded above by $O(n\log n)$, where $n=|V(G)|$. The constant implicit in the $O$-notation depends on $\lambda$, $\Delta$ and $\eta$.
@@
This follows from [9, Thm 9 and Lemma 10], [^1] The quoted results are stated for edge-based configurations rather than vertex-based, but this is inessential. which builds on earlier work, such as [10]. In order to make the correspondence we need to discuss briefly the notion of ‘pinning’. The distributions $\mu_{G,\lambda}^{(v,0)}$ and $\mu_{G,\lambda}^{(v,1)}$ are examples of distributions obtained by pinning, in this instance by pinning vertex $v$ to 0 and 1, respectively. More generally, we may consider pinning a set of vertices, say $\Lambda\subset V(G)$, to values given by $\tau:\Lambda\to\{0,1\}$, resulting in a marginal distribution $\mu_{G,\lambda}^{(\Lambda,\tau)}$ on $V(G)\setminus\Lambda$. Lemma 10 of [9] quantifies over pairs of pinnings $(\Lambda,\tau)$ and $(\Lambda,\tau^{\prime})$ where $\tau$ and $\tau^{\prime}$ differ at exactly one point. However, the hard-core distribution has a pleasant property: pinning a vertex $v$ to 0 is equivalent to removing $v$ from the graph, while pinning $v$ to 1 is equivalent to removing $v$ and all its neighbours. So, in the case of the hard-core model we need only quantify over single vertex pinnings, since we are restricting attention to graph classes which are closed under taking induced subgraphs.
From our premisses, Lemma 10 of [9] tells us that that $\mu_{G,\lambda}$ is ‘$\eta$-spectrally bounded’. For our purposes we don’t need to define this concept as we are merely going to feed this information about $\mu_{G,\lambda}$ into Theorem 9 of [9]. The remaining premiss required for this theorem is that $\mu_{G,\lambda}$ is ‘$b$-marginally bounded’. This just means that $\mathop{\mathbb{P\hbox{}}}(\sigma(v)=1)$ is bounded away from both 0 and 1 by a margin $b$. But this is exactly what is guaranteed by Corollary 3, with $b=\min\{\lambda/(1+\lambda)^{\Delta+1},1/(1+\lambda)\}$. ∎
@@ 3. $S_{1,1,t}$-free graphs (claw-free and upwards)
Suppose $\Delta\geq 3$ is a degree bound and $H$ a fixed graph. Let $G$ be an $H$-free (i.e., excluding induced copies of $H$) connected graph of maximum degree $\Delta$. Chen and Gu [9, Thm 4] show optimal mixing of Glauber dynamics for the hard-core model on $G$, for any given fugacity $\lambda>0$, when the excluded graph $H$ is the claw. In fact, optimal mixing also holds for $H=S_{1,1,t}$ with $t\geq 2$, for a rather dull reason. Define
@@
  • $A=\{\{v,w_{1}\}\}$: the set $\{v,a,b,w_{1},\ldots,w_{t}\}$ induces $S_{1,1,t}$;
  • $A=\{\{a,w_{1}\}\}$: the set $\{v,b,c,a,w_{1},\ldots,w_{t-1}\}$ induces $S_{1,1,t}$;
  • $A=\{\{v,w_{1}\},\{a,w_{1}\}\}$: the set $\{v,b,c,w_{1},\ldots,w_{t}\}$ induces $S_{1,1,t}$;
  • $A\supseteq\{\{a,w_{1}\},\{b,w_{1}\}\}$: the set $\{w_{1},a,b,w_{2},\ldots,w_{t+1}\}$ induces $S_{1,1,t}$.
@@
This is immediate from [9, Thm 4] and Corollary 6. ∎
@@ 4. $S_{1,2,t}$-free graphs (E-free and upwards)
Before heading off to larger graph classes it is instructive to go back to a precursor of [9]. Van den Berg and Brouwer [26] construct an explicit coupling for matchings in a bounded-degree graph that (in modern terms) bounds spectral independence. [^2] Although this reference is closest to our application, the general concept of ‘disagreement percolation’ appears earlier in work of van den Berg and Maes [27] and van den Berg and Steif [28]. This almost immediately gives optimal mixing of Glauber dynamics for the monomer-dimer model at all temperatures. (Of course, van den Berg and Brouwer did not have the spectral independence technology, so they were only able to deduce low-degree polynomial mixing time.) Their method extends easily to independent sets in claw-free graphs (though this was not stated in their paper). In this section, we rederive the result of Chen and Gu, as the proof will provide a base from which we may progress to large subdivided claws.
  • Let $R$ be a random ‘red’ independent set from $\mu_{G,\lambda}^{(v,1)}$.
@@
  • Consider the subgraph $G[R\cup B]$ of $G$ induced by the union of the two independent sets. Let $C\subseteq V(G)$ be the connected component of $G[R\cup B]$ containing $v$. We refer to $C$ as a cluster.
@@
Let $G$ be a connected claw-free graph, $v\in V(G)$ be a vertex, and $\lambda>0$. Then the coupling of Figure 3 is correct: $R$ is a sample from $\mu_{G,\lambda}^{(v,1)}$ and $B$ is a sample from $\mu_{G,\lambda}^{(v,0)}$.
@@
  Cluster $\textsf{Cluster BFS}(v)$  // $R$ and $B$ are random independent sets, revealed as the algorithm progresses.
@@
The claimed bound on $W_{1}$ is evidenced by the coupling just analysed. As $R\oplus B=C$, we just have to demonstrate that $|C|$ is small in expectation. Note that $G[C]$ is bipartite. It is straightforward to see that a connected bipartite claw-free graph can have no vertices of degree 3 or greater, so is either a path or a cycle. As paths and cycles are one-dimensional structures, we do not expect them to propagate far, and this is indeed the case. Imagine that we reveal the colours of vertices in a breadth-first fashion, starting with $v$ at time 0. The breadth-first search is presented as Algorithm 1. This is a little heavy for current purposes, but its generality will be useful later. As the cluster is either a path or a cycle, we have $|L_{j}|\leq 2$ for all levels $j$. Since the hard core distribution is dominated by the product of independent Bernoullli distributions (Lemma 2), on each iteration, with probability at least $p=(1/(\lambda+1))^{2\Delta}$, the breadth-first search will halt. So $\mathop{\mathbb{E\hbox{}}}|C|=\mathop{\mathbb{E\hbox{}}}[|R\oplus B|]$ is bounded above by the expectation of a random variable having a geometric distribution with success probability $(\lambda+1)^{-2\Delta}$, which is $(\lambda+1)^{2\Delta}$. ∎
@@
In light of Corollary 6, the first interesting case beyond claw free is E-free (i.e., $S_{1,2,2}$-free). Figure 4 illustrates an infinite family of E-free graphs which are not claw free. Indeed, by iterating the basic building block either horizontally or vertically, or both, we can appreciate that the family includes graphs that contain induced copies of $S_{1,1,t}$ for arbitrarily large $t$. (An instance of a fork is picked out in red in Figure 4.) So, by excluding E, we have escaped from the constraints of Corollary 6 into new, apparently interesting families of graphs.
@@
The cluster containing $v$ is bipartite, and hence, by Lemma 13 and Lozin’s decomposition theorem, is either of bounded size or is the blow-up of a path or cycle. The layers $L_{i}$ created by the breadth-first search of Algorithm 1 are of bounded size, so the depth of the search is again bounded by a geometric random variable. ∎